# Issue2Test: Generating Reproducing Test Cases from Issue Reports

### v2

@@ 1. Introduction
Issue reports, e.g., on GitHub, are commonly used to describe bugs, missing features, and other ways to improve a software project. Because addressing issues takes significant developer effort, recent work invests heavily in automated issue solving. Motivated by benchmarks of issues in popular open-source projects, such as SWE-bench (Jimenez et al., 2024), foundation models and Large Language Model (LLM)-based software engineering agents compete in their ability to successfully solve issues (Zhang et al., 2024; Roychoudhury et al., 2025; Rondon et al., 2025; Yang et al., 2024; Wang et al., 2025; Chen et al., 2024b; Tao et al., 2025; Xia et al., 2025; Pezzè et al., 2025). In addition to approaches that directly work on issues, automated program repair (Le Goues et al., 2019) tries to address bugs that manifest through a failing test. Various techniques have been explored for automated bug repair (Fan et al., 2023; Xia and Zhang, 2022; Bouzenia et al., 2025).
@@
A crucial prerequisite for successfully solving an issue or repairing a bug is a test case that reproduces the problem. Unfortunately, issue reports rarely include executable tests that reproduce the issue, making it challenging for developers and automated techniques to validate potential patches. Specifically in open-source projects, contributors often lack the expertise and familiarity to construct meaningful test cases alongside their reports. Most work on automated program repair repair assumes reproducing test cases to be given. In practice, however, developers often write test cases only after a problem was fixed (Cheng et al., 2025). Automated issue solving techniques sometimes include a step for reproducing the issue (Ruan et al., 2025), but do not explicitly focus on this important part.
@@
One approach for obtaining an issue-reproducing test case could be automated test generation. Both traditional approaches (Fraser and Arcuri, 2011; Pacheco et al., 2007) and neural methods (Watson et al., 2020; Dinella et al., 2022) attempt to reduce the time spent on manually writing tests. However, earlier techniques often fail to generate human-readable and maintainable test cases that align with project conventions and effectively validate intended program behavior (Daka et al., 2017). The emergence of LLMs has significantly enhanced automated test generation (Ziegler et al., 2022; Vaithilingam et al., 2022; Moradi Dakhel et al., 2023). LLMs can synthesize tests with human-like readability and writing style, reflecting patterns learned from extensive training on large code repositories (Ziegler et al., 2022). This capability has prompted the use of LLMs in test code generation (Bareiß et al., 2022; Nashid et al., 2023; Lemieux et al., 2023; Siddiq et al., 2023; Schäfer et al., 2023b; Chen et al., 2024a; Wang et al., 2024). While useful for creating regression tests, none of these approaches addresses the problem of creating tests that reproduce a specific issue.
@@
To address the challenge of creating issue-reproducing tests, recent work proposes test generators specifically for this purpose. Libro (Kang et al., 2023) generates tests from bug reports using LLM prompting but focuses on structured benchmarks, such as Defects4J, rather than real-world issue reproduction. Mündler et al. (Mündler et al., 2024) introduce SWT-bench-lite, a subset of SWE-bench, for evaluating issue reproduction, and modify the prompts of existing agents for this task. Yet, these agents are neither designed nor effective at generating issue-reproducing tests. Most recently, Auto-TDD (Ahmed et al., 2024) proposes a structured multi-step process, achieving a 21.7% fail-to-pass rate on SWT-bench-lite. While Auto-TDD improves over prior work, it does not explicitly enforce the creation of failing tests, which is important to successfully reproduce issues.
@@
We evaluate Issue2Test on all 276 instances of SWT-bench-lite, a subset of SWE-bench that has been extensively used for evaluating prior work (Mündler et al., 2024; Ahmed et al., 2024). The results show that our approach: i) outperforms all the existing baselines by generating tests for 91 instances (13 more than the best baseline) with the best performing LLM, ii) successfully generates tests for 20 unique issues missed by any prior work and contributes 60.4% of the union of reproduced issues by any technique, iii) imposes moderate costs of 5.21 cents per issue, which is orders of magnitude cheaper than what a human software developer would charge for performing the same task.
@@ 2.1. Running Example
Figure 1. Django-12284 issue from SWT-bench (Mündler et al., 2024).
@@ 2.3. Phase 1: Issue Comprehension
Inspired by ExecutionAgent (Bouzenia and Pradel, 2024), we employ the principle of meta-prompting which enables the LLM to dynamically retrieve structured guidelines tailored to the repository before generation. Meta-prompting serves as an adaptive knowledge retrieval mechanism, leveraging the LLM’s pre-trained knowledge to synthesize project-specific testing conventions without relying on predefined templates. For instance, we send a query to the LLM, incorporating the repository name, version, and issue description alongside instructions to generate concise guidelines for writing executable unit tests. The LLM’s response is then used in further prompts to provide project-specific context for the test generation task.
@@ 3.1.1. Benchmark
We use the SWT-bench-lite benchmark, proposed recently by Mündler et al. (Mündler et al., 2024), as the benchmark for our evaluation. SWT-bench-lite is a curated subset of the SWT-bench dataset, designed to cost-efficiently evaluate the ability of language models to generate unit tests that reproduce real-world software issues. It consists of 276 instances derived from popular GitHub repositories, each containing a GitHub issue description, a corresponding issue-resolving patch, and a set of golden reference test cases that fail on the original version but pass after the issue has been resolved. Pull requests included in SWT-bench-lite meet the following criteria: (i) they are merged into the main branch, (ii) they explicitly link to a resolved GitHub issue, and (iii) they modify at least one test file.
In addition to the full SWT-bench-lite dataset (called all issues), we also report results on those 90 instances from SWT-bench-lite that are part of SWE-bench Verified (Chowdhury et al., 2024) (called verified issues). SWE-bench Verified (Chowdhury et al., 2024) is a human-validated subset of the original SWE-bench (Jimenez et al., 2024). The validation process for SWE-bench Verified, conducted by OpenAI, involved expert annotation to ensure that each GitHub issue is solvable, clear, and linked to a valid bug-fixing patch. Each instance was manually reviewed to confirm that the issue description provides sufficient context and that the patch and test cases accurately capture the issue and its resolution.
@@ 3.1.2. LLM Empowering Issue2Test
To assess the impact of different models, we further evaluate Issue2Test with Meta Llama 3.3-70B (version meta.llama3-3-70b-instruct-v1:0 from AWS Bedrock) (Llama, 2025), an open-source model, and Claude 3.5 Sonnet (version 20241022) (Anthropic, 2025), a leading proprietary model from Anthropic.
@@ 3.1.3. Evaluation Metrics
To evaluate the effectiveness of generated test cases, we use the fail-to-pass (F→P) rate, as used in prior work (Mündler et al., 2024; Ahmed et al., 2024). Given one test generated per issue, the metric measures the proportion of generated tests that are issue-reproducing and patch-validating, i.e., they initially fail on the pre-patch code and pass after the issue-resolving patch is applied. A high F→P rate indicates that the generated tests correctly reproduce the issue and validate its resolution, making it a crucial measure of test generation quality.
@@ 3.1.4. Baselines
To evaluate Issue2Test, we compare it against several baselines that represent different approaches to test generation, following prior work by Mündler et al. (Mündler et al., 2024) and Auto-TDD (Ahmed et al., 2024). Zero-Shot Prompting (ZeroShot) uses direct LLM prompting to generate test cases based on issue descriptions. Libro (Kang et al., 2023) was the first approach specifically designed for generating bug-reproducing test cases. Code agents SWE-Agent (Yang et al., 2024) and AutoCodeRover (Zhang et al., 2024) were originally developed for issue solving, but were adapted for test generation using a prompt to create issue-reproducing unit tests (Mündler et al., 2024). SWE-Agent+ (Mündler et al., 2024) extends SWE-Agent by executing the generated tests before finalizing them, improving issue-specific test generation. Finally, Auto-TDD (Ahmed et al., 2024) is a recent approach focusing on test generation for GitHub issues similar to Issue2Test.
To compare Issue2Test with publicly available techniques, we include non-peer-reviewed entries from the SWT-Bench Lite leaderboard [^1] https://swtbench.com. This results in the inclusion of two variants of OpenHands (Wang et al., 2025), a generic-purpose coding agent. The first variant, OpenHands-Vanilla, uses the agent as released; the second variant, OpenHands-CI, augments the agent with a test‐code‐specific feedback extraction mechanism.
@@ 4. Discussion
Quality of Generated Tests. Following studies on the quality of generated tests and their potential smells (Galindo-Gutierrez et al., 2023, 2025), we manually inspect a sample of successfully generated tests to check for smells and quality issues. First, we randomly sample, up to three successfully generated tests per project (some projects have fewer than three). This results in a total of 21 tests from 8 projects. Second, we manually go through the generated tests and their corresponding execution traces. For each generated test, we look for the categories of smells described in (Galindo-Gutierrez et al., 2023), namely: (1) Act–Assert Mismatch, (2) Redundant Code, (3) Failed Setup, (4) Accessors and Constants. We only find instances of redundant code in the inspected sample. Specifically, 5/21 tests had a duplicate setup, i.e., two or more test functions instantiate the same variables with identical values and identical subsequent changes to the variables. Furthermore, 2/21 tests had a duplicate test scenario, i.e, the same test scenario is implemented in two different test cases. The percentage of these two effects is similar to the values reported in prior work (Galindo-Gutierrez et al., 2023). On the positive side, Issue2Test does not show any smell symptoms from the other three categories. One reason for the duplication may be that our prompt instructs the LLM to distribute different test scenarios and covered branches over different functions, to isolate points of failure. Future work could investigate how to guide LLMs toward tests that avoid duplication and other test smells.
Reproduction Success by Issue Category. To assess whether Issue2Test depends on the kind of issue addressed, we classified all 276 SWT-bench-lite instances according to the SWE-bench taxonomy (Jimenez et al., 2024). The full set comprises 53 bug fixes (19.2%), 11 feature requests (4.0%), no regressions (0.0%), and 212 other issues (76.8 %), the latter dominated by maintenance-oriented tags such as “help wanted”. On the 84 issues for which Issue2Test generated reproducing tests, the distribution is 17 bugs (20.2 %), 4 features (4.8 %), 0 regressions, and 63 Others (75.0 %). These correspond to per-category success rates of 32.1% for bugs (17/53), 36.4% for feature requests (4/11), and 29.7% for other issues (63/212). The maximum gap across category is 6.7%, indicating that Issue2Test performs consistently on bug fixes, feature requests, and maintenance tasks. Improving reproduction accuracy for maintenance-style (“Other”) issues may require incorporating richer contextual information, such as more detailed task descriptions, to guide the synthesis of targeted and precise test cases.
@@ 6. Related Work
Reproducible Test Case Generation. Reproducing software failures from issue reports is essential for debugging (Beller et al., 2018), regression testing (Kang et al., 2023), and automated program repair (Rondon et al., 2025). In Section 3.1.4, we discussed approaches for reproducible test case generation. Additionally, BRT Agent (Cheng et al., 2025) focused on generating tests from bug reports, specifically in an industrial setup. The reported success of BRT demonstrates that LLM-based test generation is beneficial not only for public issues but also for private internal issues at large companies such as Google.
@@
Benchmarks for Evaluating Test Generation from GitHub Issues. Several benchmarks have been developed to evaluate the effectiveness of automated test generation from GitHub issues. These benchmarks provide real-world issues, developer-written patches, and test cases, enabling systematic assessment of LLM-based test generation techniques. SWE-bench (Jimenez et al., 2024) and its derivatives such SWTBench (Mündler et al., 2024), and TDD-Bench-Verified (Ahmed et al., 2024) focus on GitHub issue resolution by pairing reported issues with their corresponding fixes and test cases. Other benchmarks, such as Defects4J (Just et al., 2014), BugsInPy (Widyasari et al., 2020), and Bugswarm (Tomassi et al., 2019) provide datasets for evaluating software testing and program repair techniques. While not specifically tailored for test generation from GitHub issues, they are still taken from real fixed issues and provide some context for bug reproduction.
@@
LLM-Based Test Generation. LLMs have been applied to test generation (Feng and Chen, 2024; Kang et al., 2023; Schäfer et al., 2023a, b; Nashid et al., 2023), with efforts to enhance quality by overcoming coverage plateaus (Lemieux et al., 2023), incorporating code-aware prompting (Ryan et al., 2024), and leveraging coverage information (Altmayer Pizzorno and Berger, 2025). LLMs have also been explored for augmenting existing tests (Alshahwan et al., 2024) refines human-written test suites to improve coverage. Our work differs by focusing not on direct code generation but on synthesizing test cases specifically for reproducing issues raised on GitHub and revealing the buggy behavior related to the issue.
@@
LLMs for GitHub Issue Resolution. LLMs are increasingly being leveraged to automate GitHub issue resolution by generating patches (Tao et al., 2025). Recent research has explored LLM-based agents that analyze issue reports, suggest code modifications, and improve software maintenance workflows. A key area of focus is automated program repair, where agents such as RepairAgent (Bouzenia et al., 2025) apply LLMs to detect and fix defects in source code. Other approaches, including SWE-Agent (Yang et al., 2024), MarsCode Agent (Liu et al., 2024), Magis (Tao et al., 2025), and AutoCodeRover (Zhang et al., 2024), tackle broader issue resolution tasks such as bug fixes, feature additions, and code enhancements. Our work, which automatically creates reproducing test cases for said issues, could help these approaches by: i) augmenting existing test suites and thus receiving more feedback about the behavior of code. ii) making the repair approaches more practical and realistic by using automatically generated test cases instead of developer-written tests, often, added after fixing the bug.
@@
Execution-Oriented Agents. ExecutionAgent (Bouzenia and Pradel, 2024) and Issue2Test both use LLMs to automate software testing tasks but differ significantly in goals and techniques. ExecutionAgent focuses on automating the setup and execution of existing test suites by inferring and running project-specific build, dependency, and test commands. It assumes tests already exist and prioritizes execution without human input. In contrast, Issue2Test generates new test cases to reproduce failures described in issue reports, operating under the assumption that no reproducing test exists. It takes issue descriptions and stack traces as input to synthesize failing test logic. This means, while ExecutionAgent focuses on command line actions such as installing dependencies, Issue2Test, in contrast, focuses on generating test code. Subsequently, meta-prompting in ExecutionAgent helps infer commands and actions tied to the repository structure and languages while Issue2Test uses it to reflect project-specific test idioms such as naming, fixtures, and assertions. Though the principle of meta-prompting is the same, Issue2Test adapts its own meta-prompting for the target task. Furthermore, Issue2Test uses a similarity-based approach to find relevant test and code files while ExecutionAgent just relies on prompting and meta-promting to find relevant files at the start.

### v3

@@ 1. Introduction
Issue reports, e.g., on GitHub, are commonly used to describe bugs, missing features, and other ways to improve a software project. Because addressing issues takes significant developer effort, recent work invests heavily in automated issue solving. Motivated by benchmarks of issues in popular open-source projects, such as SWE-bench (swebench), foundation models and Large Language Model (LLM)-based software engineering agents compete in their ability to successfully solve issues (autocoderover; roy:ai-software-engineer:arxiv25; rondon:agent-based-repair-google:arxiv25; sweagent:arxiv24; openhands:iclr25; coder:arxiv24; magis:neurips25; agentless:arxiv24; pezz:roadmap-se:tosem25). In addition to approaches that directly work on issues, automated program repair (cacm2019-program-repair) tries to address bugs that manifest through a failing test. Various techniques have been explored for automated bug repair (fan:code-repair-with-llm:icse23; xia:alpharepair:fse22; repairagent:icse25).
@@
A crucial prerequisite for successfully solving an issue or repairing a bug is a test case that reproduces the problem. Unfortunately, issue reports rarely include executable tests that reproduce the issue, making it challenging for developers and automated techniques to validate potential patches. Specifically in open-source projects, contributors often lack the expertise and familiarity to construct meaningful test cases alongside their reports. Most work on automated program repair repair assumes reproducing test cases to be given. In practice, however, developers often write test cases only after a problem was fixed (cheng:agentic-bug-reproduction:arxiv25). Automated issue solving techniques sometimes include a step for reproducing the issue (specrover:arxiv24), but do not explicitly focus on this important part.
@@
One approach for obtaining an issue-reproducing test case could be automated test generation. Both traditional approaches (evosuite:fse11; pacheco:randoop:icse07) and neural methods (atlas:test-assert-generation:icse20; toga:icse22) attempt to reduce the time spent on manually writing tests. However, earlier techniques often fail to generate human-readable and maintainable test cases that align with project conventions and effectively validate intended program behavior (daka:unit-test-readability:issta17). The emergence of LLMs has significantly enhanced automated test generation (ziegler:copilot-productivity:maps:22; vaithilingam:copilot:chi22; dakhel2022github). LLMs can synthesize tests with human-like readability and writing style, reflecting patterns learned from extensive training on large code repositories (ziegler:copilot-productivity:maps:22). This capability has prompted the use of LLMs in test code generation (bareiss:test-generation-llm:arxiv22; cedar; lemieux:codamosa:icse23; siddiq:llm-unit-test:arxiv23; schafer:unit-test-generation-using-llm:tse23; xie:chatunitest:arxiv23; software-testing-with-llm-survey). While useful for creating regression tests, none of these approaches addresses the problem of creating tests that reproduce a specific issue.
@@
To address the challenge of creating issue-reproducing tests, recent work proposes test generators specifically for this purpose. Libro (libro:icse23) generates tests from bug reports using LLM prompting but focuses on structured benchmarks, such as Defects4J, rather than real-world issue reproduction. Mündler et al. (niels:bug-fixes-with-code-agents:nips24) introduce SWT-bench-lite, a subset of SWE-bench, for evaluating issue reproduction, and modify the prompts of existing agents for this task. Yet, these agents are neither designed nor effective at generating issue-reproducing tests. Most recently, Auto-TDD (autotdd:arxiv24) proposes a structured multi-step process, achieving a 21.7% fail-to-pass rate on SWT-bench-lite. While Auto-TDD improves over prior work, it does not explicitly enforce the creation of failing tests, which is important to successfully reproduce issues.
@@
We evaluate Issue2Test on all 276 instances of SWT-bench-lite, a subset of SWE-bench that has been extensively used for evaluating prior work (niels:bug-fixes-with-code-agents:nips24; autotdd:arxiv24). The results show that our approach: i) outperforms all the existing baselines by generating tests for 91 instances (13 more than the best baseline) with the best performing LLM, ii) successfully generates tests for 20 unique issues missed by any prior work and contributes 60.4% of the union of reproduced issues by any technique, iii) imposes moderate costs of 5.21 cents per issue, which is orders of magnitude cheaper than what a human software developer would charge for performing the same task.
@@ 2.1. Running Example
Figure 1. Django-12284 issue from SWT-bench (niels:bug-fixes-with-code-agents:nips24).
@@ 2.3. Phase 1: Issue Comprehension
Inspired by ExecutionAgent (agent-test-execution:arxiv24), we employ the principle of meta-prompting which enables the LLM to dynamically retrieve structured guidelines tailored to the repository before generation. Meta-prompting serves as an adaptive knowledge retrieval mechanism, leveraging the LLM’s pre-trained knowledge to synthesize project-specific testing conventions without relying on predefined templates. For instance, we send a query to the LLM, incorporating the repository name, version, and issue description alongside instructions to generate concise guidelines for writing executable unit tests. The LLM’s response is then used in further prompts to provide project-specific context for the test generation task.
@@ 3.1.1. Benchmark
We use the SWT-bench-lite benchmark, proposed recently by Mündler et al. (niels:bug-fixes-with-code-agents:nips24), as the benchmark for our evaluation. SWT-bench-lite is a curated subset of the SWT-bench dataset, designed to cost-efficiently evaluate the ability of language models to generate unit tests that reproduce real-world software issues. It consists of 276 instances derived from popular GitHub repositories, each containing a GitHub issue description, a corresponding issue-resolving patch, and a set of golden reference test cases that fail on the original version but pass after the issue has been resolved. Pull requests included in SWT-bench-lite meet the following criteria: (i) they are merged into the main branch, (ii) they explicitly link to a resolved GitHub issue, and (iii) they modify at least one test file.
In addition to the full SWT-bench-lite dataset (called all issues), we also report results on those 90 instances from SWT-bench-lite that are part of SWE-bench Verified (swebench-verified) (called verified issues). SWE-bench Verified (swebench-verified) is a human-validated subset of the original SWE-bench (swebench). The validation process for SWE-bench Verified, conducted by OpenAI, involved expert annotation to ensure that each GitHub issue is solvable, clear, and linked to a valid bug-fixing patch. Each instance was manually reviewed to confirm that the issue description provides sufficient context and that the patch and test cases accurately capture the issue and its resolution.
@@ 3.1.2. LLM Empowering Issue2Test
To assess the impact of different models, we further evaluate Issue2Test with Meta Llama 3.3-70B (version meta.llama3-3-70b-instruct-v1:0 from AWS Bedrock) (llama), an open-source model, and Claude 3.5 Sonnet (version 20241022) (claude), a leading proprietary model from Anthropic.
@@ 3.1.3. Evaluation Metrics
To evaluate the effectiveness of generated test cases, we use the fail-to-pass (F→P) rate, as used in prior work (niels:bug-fixes-with-code-agents:nips24; autotdd:arxiv24). Given one test generated per issue, the metric measures the proportion of generated tests that are issue-reproducing and patch-validating, i.e., they initially fail on the pre-patch code and pass after the issue-resolving patch is applied. A high F→P rate indicates that the generated tests correctly reproduce the issue and validate its resolution, making it a crucial measure of test generation quality.
@@ 3.1.4. Baselines
To evaluate Issue2Test, we compare it against several baselines that represent different approaches to test generation, following prior work by Mündler et al. (niels:bug-fixes-with-code-agents:nips24) and Auto-TDD (autotdd:arxiv24). Zero-Shot Prompting (ZeroShot) uses direct LLM prompting to generate test cases based on issue descriptions. Libro (libro:icse23) was the first approach specifically designed for generating bug-reproducing test cases. Code agents SWE-Agent (sweagent:arxiv24) and AutoCodeRover (autocoderover) were originally developed for issue solving, but were adapted for test generation using a prompt to create issue-reproducing unit tests (niels:bug-fixes-with-code-agents:nips24). SWE-Agent+ (niels:bug-fixes-with-code-agents:nips24) extends SWE-Agent by executing the generated tests before finalizing them, improving issue-specific test generation. Finally, Auto-TDD (autotdd:arxiv24) is a recent approach focusing on test generation for GitHub issues similar to Issue2Test.
To compare Issue2Test with publicly available techniques, we include non-peer-reviewed entries from the SWT-Bench Lite leaderboard [^1] https://swtbench.com. This results in the inclusion of two variants of OpenHands (openhands:iclr25), a generic-purpose coding agent. The first variant, OpenHands-Vanilla, uses the agent as released; the second variant, OpenHands-CI, augments the agent with a test‐code‐specific feedback extraction mechanism.
@@ 4. Discussion
Quality of Generated Tests. Following studies on the quality of generated tests and their potential smells (10336341; galindo2025increasing), we manually inspect a sample of successfully generated tests to check for smells and quality issues. First, we randomly sample, up to three successfully generated tests per project (some projects have fewer than three). This results in a total of 21 tests from 8 projects. Second, we manually go through the generated tests and their corresponding execution traces. For each generated test, we look for the categories of smells described in (10336341), namely: (1) Act–Assert Mismatch, (2) Redundant Code, (3) Failed Setup, (4) Accessors and Constants. We only find instances of redundant code in the inspected sample. Specifically, 5/21 tests had a duplicate setup, i.e., two or more test functions instantiate the same variables with identical values and identical subsequent changes to the variables. Furthermore, 2/21 tests had a duplicate test scenario, i.e, the same test scenario is implemented in two different test cases. The percentage of these two effects is similar to the values reported in prior work (10336341). On the positive side, Issue2Test does not show any smell symptoms from the other three categories. One reason for the duplication may be that our prompt instructs the LLM to distribute different test scenarios and covered branches over different functions, to isolate points of failure. Future work could investigate how to guide LLMs toward tests that avoid duplication and other test smells.
Reproduction Success by Issue Category. To assess whether Issue2Test depends on the kind of issue addressed, we classified all 276 SWT-bench-lite instances according to the SWE-bench taxonomy (swebench). The full set comprises 53 bug fixes (19.2%), 11 feature requests (4.0%), no regressions (0.0%), and 212 other issues (76.8 %), the latter dominated by maintenance-oriented tags such as “help wanted”. On the 84 issues for which Issue2Test generated reproducing tests, the distribution is 17 bugs (20.2 %), 4 features (4.8 %), 0 regressions, and 63 Others (75.0 %). These correspond to per-category success rates of 32.1% for bugs (17/53), 36.4% for feature requests (4/11), and 29.7% for other issues (63/212). The maximum gap across category is 6.7%, indicating that Issue2Test performs consistently on bug fixes, feature requests, and maintenance tasks. Improving reproduction accuracy for maintenance-style (“Other”) issues may require incorporating richer contextual information, such as more detailed task descriptions, to guide the synthesis of targeted and precise test cases.
@@ 6. Related Work
Reproducible Test Case Generation. Reproducing software failures from issue reports is essential for debugging (beller:debugging:icse18), regression testing (libro:icse23), and automated program repair (rondon:agent-based-repair-google:arxiv25). In Section 3.1.4, we discussed approaches for reproducible test case generation. Additionally, BRT Agent (cheng:agentic-bug-reproduction:arxiv25) focused on generating tests from bug reports, specifically in an industrial setup. The reported success of BRT demonstrates that LLM-based test generation is beneficial not only for public issues but also for private internal issues at large companies such as Google.
@@
Benchmarks for Evaluating Test Generation from GitHub Issues. Several benchmarks have been developed to evaluate the effectiveness of automated test generation from GitHub issues. These benchmarks provide real-world issues, developer-written patches, and test cases, enabling systematic assessment of LLM-based test generation techniques. SWE-bench (swebench) and its derivatives such SWTBench (niels:bug-fixes-with-code-agents:nips24), and TDD-Bench-Verified (autotdd:arxiv24) focus on GitHub issue resolution by pairing reported issues with their corresponding fixes and test cases. Other benchmarks, such as Defects4J (defects4j), BugsInPy (bugsinpy), and Bugswarm (bugswarm) provide datasets for evaluating software testing and program repair techniques. While not specifically tailored for test generation from GitHub issues, they are still taken from real fixed issues and provide some context for bug reproduction.
@@
LLM-Based Test Generation. LLMs have been applied to test generation (feng:bug-replay-llm:icse24; libro:icse23; schafer:test-generation-using-llm:arxiv23; schafer:unit-test-generation-using-llm:tse23; cedar), with efforts to enhance quality by overcoming coverage plateaus (lemieux:codamosa:icse23), incorporating code-aware prompting (symprompt), and leveraging coverage information (coverup:arxiv24). LLMs have also been explored for augmenting existing tests (test-improvement-using-llms-meta) refines human-written test suites to improve coverage. Our work differs by focusing not on direct code generation but on synthesizing test cases specifically for reproducing issues raised on GitHub and revealing the buggy behavior related to the issue.
@@
LLMs for GitHub Issue Resolution. LLMs are increasingly being leveraged to automate GitHub issue resolution by generating patches (magis:neurips25). Recent research has explored LLM-based agents that analyze issue reports, suggest code modifications, and improve software maintenance workflows. A key area of focus is automated program repair, where agents such as RepairAgent (repairagent:icse25) apply LLMs to detect and fix defects in source code. Other approaches, including SWE-Agent (sweagent:arxiv24), MarsCode Agent (marscode-agent:arxiv24), Magis (magis:neurips25), and AutoCodeRover (autocoderover), tackle broader issue resolution tasks such as bug fixes, feature additions, and code enhancements. Our work, which automatically creates reproducing test cases for said issues, could help these approaches by: i) augmenting existing test suites and thus receiving more feedback about the behavior of code. ii) making the repair approaches more practical and realistic by using automatically generated test cases instead of developer-written tests, often, added after fixing the bug.
@@
Execution-Oriented Agents. ExecutionAgent (agent-test-execution:arxiv24) and Issue2Test both use LLMs to automate software testing tasks but differ significantly in goals and techniques. ExecutionAgent focuses on automating the setup and execution of existing test suites by inferring and running project-specific build, dependency, and test commands. It assumes tests already exist and prioritizes execution without human input. In contrast, Issue2Test generates new test cases to reproduce failures described in issue reports, operating under the assumption that no reproducing test exists. It takes issue descriptions and stack traces as input to synthesize failing test logic. This means, while ExecutionAgent focuses on command line actions such as installing dependencies, Issue2Test, in contrast, focuses on generating test code. Subsequently, meta-prompting in ExecutionAgent helps infer commands and actions tied to the repository structure and languages while Issue2Test uses it to reflect project-specific test idioms such as naming, fixtures, and assertions. Though the principle of meta-prompting is the same, Issue2Test adapts its own meta-prompting for the target task. Furthermore, Issue2Test uses a similarity-based approach to find relevant test and code files while ExecutionAgent just relies on prompting and meta-promting to find relevant files at the start.